OpenStudy (marieokay12345seven):

Solve the system 2x + 2y = -6 and 3x - 2y = 11 by using graph paper or graphing technology. What is the solution to the system? (2 points)

OpenStudy (marieokay12345seven):

yup! (-1, -7) (1, -4) (2, -1) (3, -2)

OpenStudy (supersmart1001):

If you have a graphing calculator rearrange the formula for y=mx + b and plug into the calculator same deal look for the intersection. You could also solve like this 3/2(2x+3y = 3) is 3x + (9/2)y = 9/2 then subtract second equation from first to get 3x + (9/2)y = 9/2 -3x - 2y = 11 becomes 13/2y = -13/2 y = -1 3(x) - 2(-1) = 11 so x = 3 So you answer should come out to___

OpenStudy (marieokay12345seven):

(3, -2) ???

OpenStudy (supersmart1001):

thats what i got ask the moderator for extra help just in case

OpenStudy (marieokay12345seven):

@.Sam. I think you're the moderator! Is the answer correct?

OpenStudy (supersmart1001):

#AFK

OpenStudy (marieokay12345seven):

@phi ?